# Counterfactual Explanations with Probabilistic Guarantees on their Robustness to Model Change

### v1

@@ C.2 Proof
This result, $t>0.5$ proves the validity of Lemma 3, which in turn validates proves Theorem 2, because the highest attainable value of CDF is at $1-\alpha$, where $\alpha>0.5$
Appendix D Background on Bernoulli and Beta Distributions
D.1 Bernoulli Distribution
$P(X=x)=p^{x}(1-p)^{1-x},\quad x\in\{0,1\}$ (21)
where $p$ is the probability of success (i.e., $X=1$).
D.2 Beta Distribution
$f(x;a,b)=\frac{x^{a-1}(1-x)^{b-1}}{B(a,b)},\quad 0\leq x\leq 1$ (22)
where $B(\alpha,\beta)$ is the Beta function.
D.3 Conjugate Relationship and Conjugate Priors
  • We then observe a series of coin flips (Bernoulli trials).
Mathematically, this relationship is expressed as:
[…]
Appendix E Examining BetaRCE parameters in detail
@@
Figure 5: Beta distribution with varying priors.
It’s evident that increasing the $a$ parameter skews the distribution to the right, while altering the $b$ parameter skews it to the left. Therefore, adding $k$ to $a$ identifies the most optimistic (positively skewed) distribution obtainable with the given parameters. Consequently, this facilitates the calculation of the most optimistic lower bound that can be attained: $\delta_{max}$. The proof for this statement is in Sec. C.
@@
Figure 6: Max achievable $\delta$ values based on varying parameters.
@@
Appendix F Experimental setup
@@ F.1 General
F.2 Datasets
Below, we present basic information about the datasets used in our study.
@@ F.2 Datasets
The preprocessing of these datasets involved dropping rows containing missing values and performing min-max normalization.
F.3 Hyperparameters of the Baselines
@@
For all visualisations, we selected the hyperparameter setting that gave the highest empirical robustness to ensure a fair comparison. The only exception is the post-hoc method RobX , because for a post-hoc method it is also important to assess the distance to the base counterfactual. Therefore, to show different aspects of RobX , we included two different settings in all comparisons: one that excels in empirical robustness and one that achieves a good distance to the base CFE.
F.4 Models
In our experiments, we employ two models as the underlying black-boxes: a neural network (NN) and LightGBM. The NN and LightGBM were implemented using the PyTorch and lightgbm Python libraries, respectively. Validation sets were used for early stopping in the NN and as the evaluation set for LightGBM. Below, we provide more details on their specifications:
F.4.1 Neural Network
@@ F.4.2 LightGBM
Appendix G BetaRCE intrinsic analysis
@@
G.1 Credible intervals for robustness
@@
Figure 8: With GrowingSpheres as a base counterfactual explanation.
G.2 Exploring the impact of the confidence parameter
@@
Our first analysis juxtaposes $\alpha$ with Empirical Robustness across three $\delta$ (Fig. 9).
Figure 9: Empirical robustness as a function of $\alpha$ and $\delta$
@@
The subsequent visualization illustrates this enhanced security with higher $\alpha$ values, as indicated by the greater distance between the blue line and the red line, representing the lower bound of the credible interval:
Figure 10: The relationship of empirical robustness and $\alpha$ confidence and its placement credible intervals
G.3 Exploring the impact of the number of estimators
In this section, we examine how varying the number of estimators, denoted as $k$, affects the performance of BetaRCE . As depicted in the Fig. 11 below, increasing $k$ results in narrower credible intervals, indicating a higher level of confidence in the robustness range.
Figure 11: The impact of parameter $k$ on the average credible interval width
This outcome is anticipated because a higher value of $k$ allows for more combinations of parameters $a$ and $b$ to form the Beta distribution. Consequently, the distribution becomes more flexible, enabling a better fit to the empirical distribution.
The next plot (Fig. 12) illustrates the relationship of the $k$ parameter and the Empirical Robustness.
Figure 12: The impact of parameter $k$ on the average empirical robustness
As observed, the average Empirical Robustness does not seem to highly depend on the number of estimators.
From these empirical experiments, the conclusion that $k$ increases the Empirical Robustness cannot be drawn. Therefore, our recommendation is to use the lowest possible $k$ which allows for realizing desired $\delta$, using introduced for that purpose equation (Eq. 7).
G.4 Investigating the generalization capabilities across different experiment types
@@
The results of these experiments are presented in three tables (Tables 45 and 6). Each table shows the results for a different base CFE method, averaged across four datasets, with $\delta=0.9$ and $\alpha=0.95$.
@@
Table 6: Empirical Robustness of BetaRCE with FACE as the base CFE generation method. The results are averaged over all datasets.
Generalization
Original Architecture Bootstrap Seed
Architecture $0.974\pm 0.003$ $0.904\pm 0.006$ $0.937\pm 0.005$
Bootstrap $0.969\pm 0.004$ $0.914\pm 0.006$ $0.962\pm 0.004$
Seed $0.962\pm 0.004$ $0.891\pm 0.006$ $0.925\pm 0.005$
Appendix H Comparative Analysis
In this section, we present the comprehensive results from the comparative study detailed in Sec. 4.4. Tab. LABEL:app:tab:nn-gs provides the extended version of Tab. 2, including results across all four datasets, with GrowingSpheres used as the base CFE method. Next, in Tab. 8, we present the results for when the Dice generates base CFEs. Finally, in Tab. 9, we include the results for the scenario where LightGBM is the underlying black-box model, and GrowingSpheres is employed as the base CFE generation method.
@@ Appendix H Comparative Analysis
Table 7: Comparative study results when LightGBM is used as a black-box model. RobX   and BetaRCE were using GrowingSpheres as the base counterfactual explainer. Parameters used in a given method are listed next to this method’s name; for RobX these are $\tau$ and variance, while for BetaRCE $\delta$ ($\alpha=0.95$). The values in each cell represent the mean $\pm$ standard error. The column Type sorts the methods by categories. The abbreviations Btsr and Arch used next to BetaRCE in the Type column stand for Bootstrap and Architecture, respectively.
@@
Table 8: Comparative study results. RobX   and BetaRCE were using Dice as the base counterfactual explainer. Parameters used in a given method are listed next to this method’s name; for RobX these are $\tau$ and variance, while for BetaRCE $\delta$ ($\alpha=0.9$). The values in each cell represent the mean $\pm$ standard error. The column Type sorts the methods by categories. The abbreviations Btsr and Arch used next to BetaRCE in the Type column stand for Bootstrap and Architecture, respectively.

### v3

@@ C.2 Proof
The result of Lemma 4, $t>0.5$, finishes the proof for Lemma 3, which in turn proves Theorem 2, because the highest attainable value of CDF is at $1-\alpha$, where $\alpha>0.5$
Appendix D Examining BetaRCE parameters in detail
@@
Figure 4: Some priors of Beta distribution
Increasing the $a$ parameter skews the distribution to the right, while altering the $b$ parameter skews it to the left. Therefore, adding $k$ to $a$ identifies the most optimistic (positively skewed) distribution obtainable with the given parameters. Consequently, this facilitates the calculation of the most optimistic lower bound that can be attained: $\delta_{max}$. The proof for this statement is in Sec. C.
@@
Figure 5: Max achievable $\delta$ values as a function of $k$ and $\alpha$.
@@
Appendix E Experimental setup
@@ E.1 General
E.2 Datasets
Below, we provide information about the datasets used in our study.
@@ E.2 Datasets
Car eval 1728 6 2.34
Rice 3810 7 1.34
Preprocessing of these datasets involved dropping rows containing missing values and performing min-max normalization.
E.3 Hyperparameters of the Baselines
@@
For all visualizations, we chose the hyperparameter configuration that provided the highest empirical robustness to ensure a fair comparison. The only exception is the post-hoc method RobX , as it is also crucial to evaluate the distance to the base counterfactual in such methods. Therefore, to highlight different aspects of RobX , we included two distinct settings in all comparisons: one optimized for empirical robustness and another that strikes a balance with a good distance to the base CFE.
E.4 Models
In our experiments, we utilize three models as the core black-boxes: a neural network (NN), LightGBM, and logistic regression (LR). These models are implemented using torch, lightgbm, and sklearn, respectively. The validation sets were employed for early stopping in the NN and as the evaluation set for LightGBM. Detailed specifications are provided below:
E.4.1 Neural Network
@@ E.4.2 LightGBM
E.4.3 Logistic Regression
Parameter Fixed hparams Hparams to Vary
solver lbfgs lbfgs, newton-cg, sag
penalty l2 l2, none
max_iter 100 50 - 200
C 1 0.1 - 1.0
seed 42
Appendix F BetaRCE intrinsic analysis
@@
F.1 Credible intervals for robustness
@@
Figure 6: With GrowingSpheres as a base counterfactual explanation.
@@
F.2 Exploring the impact of the confidence parameter
@@
Our first analysis juxtaposes $\alpha$ with Empirical Robustness across three $\delta$ (Fig. 8(a)).
@@
The subsequent visualization (Fig 8(b)) illustrates this enhanced security with higher $\alpha$ values, as indicated by the greater distance between the blue line and the red line, representing the lower bound of the credible interval:
(a) Empirical robustness as a function of $\alpha$ and $\delta$
(b) The relationship of empirical robustness and $\alpha$ confidence and its placement credible intervals
F.3 Exploring the impact of the number of estimators
In this section, we examine how varying the number of estimators, denoted as $k$, affects the performance of BetaRCE . As depicted in Fig. 9, increasing $k$ results in narrower credible intervals, indicating a higher level of confidence in the robustness range. Additionally, the relationship between $k$ and the Empirical Robustness is illustrated.
Figure 9: The impact of parameter $k$ on the performance of BetaRCE . The subfigure on the left side shows how $k$ affects the credible interval width, while the one on the right side illustrates the relationship between $k$ and Empirical Robustness.
Increasing $k$ allows for more combinations of parameters $a$ and $b$ to form the Beta distribution, making the distribution more flexible and better suited to the empirical data. Consequently, this leads to narrower credible intervals.
However, as observed in the right pane of Fig. 9, the average Empirical Robustness does not appear to strongly depend on $k$. From these empirical experiments, the conclusion that increasing $k$ enhances Empirical Robustness cannot be drawn. Therefore, our recommendation is to use the lowest possible $k$ that achieves the desired $\delta$, as determined by Eq. 7.
F.4 Investigating the generalization capabilities across different experiment types
@@
The results of these experiments are presented in two tables (Tables 3 and 4). Each table shows the results for a different base CFE method, averaged across four datasets, with $\delta=0.9$ and $\alpha=0.95$.
@@
Appendix G Computational Complexity
The computational complexity of BetaRCE is determined primarily by two factors: (1) the complexity of the chosen optimization algorithm and (2) the number of estimators ($k$) used for bootstrap robustness verification. The latter directly affects the complexity of evaluating the objective function’s constraints, as each evaluation requires querying $k$ estimators. Consequently, $k$ inference calls are introduced as a constant multiplier to the overall complexity.
The optimization algorithm employed to solve Eq. 6 plays the most significant role in determining BetaRCE’s complexity. In this paper, we utilize GrowingSpheres for optimization. While this algorithm does not offer theoretical guarantees regarding the complexity of finding an optimal or $\epsilon$-optimal solution, its complexity is heavily influenced by $\eta$ and $n$ hyperparameters.
First, $\eta$ controls the granularity of the iterative expansion of the sphere’s perimeter. Second, $n$ specifies the number of samples evaluated on the perimeter during each iteration. When combined with BetaRCE, each step of GrowingSpheres requires $n\cdot k+1$ model evaluations. Here, $k$ estimators perform inference on each of the $n$ samples to compute the robustness term (Eq. 5), while the additional $+1$ accounts for the evaluation of the validity term (Eq. 4).
Figure 10: Robust counterfactual generation time as a function of $\delta$.
Appendix H Step-by-step visualization of BetaRCE algorithm
In this section, we provide an intuitive visualization of how BetaRCE works when integrated with GrowingSpheres . Fig. 11 illustrates eight consecutive steps of the BetaRCE algorithm for a simple example.
First, a base counterfactual explanation is generated for a given input (see Fig. 11(a)). Next, we move to the warmup stage of the GrowingSpheres search algorithm, as described in Alg. 2 (lines 2-6). In particular, first, in Fig. 11(b) five candidates are sampled from a large sphere (line 2), second, in Fig. 11(c) robustness and validity are evaluated (line 3). Since there were both valid and robust examples in the sphere, its radius is halved (line 4). Next, Fig. 11(d) and Fig. 11(e) show similar steps for a sphere with smaller radius (line 5), but this time, no valid and robust examples were found. Thus, the warmup stage of GrowingSpheres is over.
In the next figures, we show an iteration of the search procedure (lines 7-13), where candidates are sampled from a region between increasing lower and upper radius. Fig. 11(f) illustrates the sampling of candidates (line 8) and Fig. 11(g) shows the candidates being evaluated (line 9). Since two robust and valid examples were found, the algorithm is terminated (lines 9-13) and the closest counterfactual (Fig. 11(h)) is returned as the robust counterfactual explanation (line 14).
Figure 11: A visualization of BetaRCE with GrowingSpheres search algorithm.
Figure 12: A plot of the correlation between distance to base and other metrics.
Appendix I The relationship between Distance to Base and other metrics
In this section, we briefly share the empirical observation that the Distance to Base metric is negatively correlated with the degradation of CFE performance metrics, focusing on different properties of CFE. Hence, we believe that post-hoc methods which introduce smaller changes to the original CFE (in terms of Distance to Base) are generally better at preserving its properties. Fig. 12 illustrates how proximity metrics (both L1 and L2) deteriorate on average as Distance to Base increases. We also note that this deterioration is significantly less prominent with the Plausibility metric.
To support these findings, we calculated correlation coefficients in Tab. 5.
Table 5: Correlation between distance to base and other metrics. All p-values are lesser than 0.001.
Metric Correlation type rho
Proximity L1 Pearson -0.31
Proximity L1 Spearman -0.70
Proximity L2 Pearson -0.31
Proximity L2 Spearman -0.73
Plausibility Pearson -0.31
Plausibility Spearman -0.02
Average Pearson -0.31
Average Spearman -0.48
Appendix J Background on Bernoulli and Beta Distributions
J.1 Bernoulli Distribution
$P(X=x)=p^{x}(1-p)^{1-x},\quad x\in\{0,1\}$ (25)
where $p$ is the probability of success (i.e., $X=1$).
J.2 Beta Distribution
$f(x;a,b)=\frac{x^{a-1}(1-x)^{b-1}}{B(a,b)},\quad 0\leq x\leq 1$ (26)
where $B(\alpha,\beta)$ is the Beta function.
J.3 Conjugate Relationship and Conjugate Priors
  • We then observe a series of coin flips (Bernoulli trials).
Mathematically, this relationship is expressed as:
[…]
Appendix K Comparative Analysis
@@ Appendix I The relationship between Distance to Base and other metrics
In this section, we present the comprehensive results from the comparative study detailed in Sec. 4.4. Parameters used in a given method are listed next to this method’s name; for RobX these are $\tau$ and variance, while for BetaRCE $\delta$ ($\alpha=0.9$). The values in each cell represent the mean $\pm$ standard error. The column Type sorts the methods by categories. The abbreviations Btsr and Arch used next to BetaRCE in the Type column stand for Bootstrap and Architecture, respectively. Below present results for three base models: neural network (NN), LightGBM, logistic regression (LR).
  • NN: Tab. 6 provides the extended version of Tab. 1, including results across all datasets, with GrowingSpheres used as the base CFE method. Additionally, in Tab. 7, we present the results for when Dice generates base CFEs.
  • LightGBM: Tab. 8 contains the results for the scenario where LightGBM is the underlying black-box model, and GrowingSpheres is employed as the base CFE generation method.
  • LR: Tab. 9 similarly provides results for GrowingSpheres as a base CFE generation method, but with logistic regressor employed as a black-box model.
Table 6: Comparative study results when the underlying model is a neural network and both RobX   and BetaRCE are using GrowingSpheres as the base counterfactual explainer
@@
BetaRCE Arch BetaRCE (0.8) 0.488 $\pm$ 0.013 0.870 $\pm$ 0.013 $0.382\pm 0.005$ 0.372 $\pm$ 0.004 0.966 $\pm$ 0.004 - -
BetaRCE (0.9) 0.607 $\pm$ 0.014 0.953 $\pm$ 0.013 $0.420\pm 0.006$ 0.378 $\pm$ 0.004 0.975 $\pm$ 0.004 - -
BetaRCE Btsr BetaRCE (0.8) 0.445 $\pm$ 0.006 0.840 $\pm$ 0.008 $0.359\pm 0.003$ 0.359 $\pm$ 0.002 - 0.903 $\pm$ 0.006 -
BetaRCE (0.9) 0.583 $\pm$ 0.007 0.949 $\pm$ 0.008 $0.407\pm 0.003$ 0.369 $\pm$ 0.002 - 0.937 $\pm$ 0.005 -
BetaRCE Seed BetaRCE (0.8) 0.247 $\pm$ 0.006 0.813 $\pm$ 0.008 $0.346\pm 0.003$ 0.350 $\pm$ 0.002 - - 0.871 $\pm$ 0.006
BetaRCE (0.9) 0.315 $\pm$ 0.006 0.862 $\pm$ 0.008 $0.367\pm 0.003$ 0.353 $\pm$ 0.002 - - 0.902 $\pm$ 0.006
BetaRCE Arch BetaRCE (0.8) 1.538 $\pm$ 0.049 2.912 $\pm$ 0.053 $0.749\pm 0.014$ 0.802 $\pm$ 0.011 0.904 $\pm$ 0.006 - -
BetaRCE (0.9) 1.697 $\pm$ 0.031 2.927 $\pm$ 0.036 $0.753\pm 0.009$ 0.783 $\pm$ 0.007 0.935 $\pm$ 0.005 - -
BetaRCE Btsr BetaRCE (0.8) 2.288 $\pm$ 0.041 3.451 $\pm$ 0.044 $0.889\pm 0.011$ 0.859 $\pm$ 0.008 - 0.833 $\pm$ 0.007 -
BetaRCE (0.9) 3.547 $\pm$ 0.071 4.501 $\pm$ 0.073 $1.156\pm 0.019$ 1.044 $\pm$ 0.015 - 0.880 $\pm$ 0.006 -
BetaRCE Seed BetaRCE (0.8) 1.420 $\pm$ 0.021 2.526 $\pm$ 0.028 $0.653\pm 0.007$ 0.726 $\pm$ 0.004 - - 0.826 $\pm$ 0.007
BetaRCE (0.9) 1.927 $\pm$ 0.030 2.906 $\pm$ 0.035 $0.750\pm 0.009$ 0.776 $\pm$ 0.006 - - 0.902 $\pm$ 0.006
Wine Standard CFEs Dice - 0.888 $\pm$ 0.001 0.549 $\pm$ 0.001 0.413 $\pm$ 0.001 0.934 $\pm$ 0.002 0.839 $\pm$ 0.003 0.873 $\pm$ 0.003
GrowingSpheres - 0.474 $\pm$ 0.001 0.174 $\pm$ 0.001 0.211 $\pm$ 0.001 0.877 $\pm$ 0.003 0.837 $\pm$ 0.003 0.877 $\pm$ 0.003
FACE - 0.54 $\pm$ 0.001 0.216 $\pm$ 0.001 0.131 $\pm$ 0.001 0.78 $\pm$ 0.003 0.747 $\pm$ 0.003 0.783 $\pm$ 0.003
Robust end-to-end RBR - 0.508 $\pm$ 0.001 0.199 $\pm$ 0.001, 0.176 $\pm$ 0.001 0.749 $\pm$ 0.002 0.73 $\pm$ 0.002 0.764 $\pm$ 0.002
ROAR - 15.768 $\pm$ 0.001 5.607 $\pm$ 0.001 5.26 $\pm$ 0.001 0.734 $\pm$ 0.002 0.755 $\pm$ 0.002 0.727 $\pm$ 0.002
Robust post-hoc RobX (0.5,0.01) 0.564 $\pm$ 0.001 0.674 $\pm$ 0.001 0.293 $\pm$ 0.001 0.159 $\pm$ 0.001 1.000 $\pm$ 0.001 0.979 $\pm$ 0.003 0.997 $\pm$ 0.001
RobX (0.5,0.1) 1.318 $\pm$ 0.002 1.378 $\pm$ 0.002 0.593 $\pm$ 0.001 0.236 $\pm$ 0.001 1.000 $\pm$ 0.001 0.970 $\pm$ 0.004 1.000 $\pm$ 0.001
Arch BetaRCE(0.8) 0.330 $\pm$ 0.005 0.525 $\pm$ 0.006 0.192 $\pm$ 0.002 0.219 $\pm$ 0.002 0.918 $\pm$ 0.005 - -
BetaRCE(0.9) 0.471 $\pm$ 0.006 0.647 $\pm$ 0.007 0.237 $\pm$ 0.003 0.242 $\pm$ 0.002 0.959 $\pm$ 0.004 - -
Btsp BetaRCE(0.8) 0.335 $\pm$ 0.005 0.513 $\pm$ 0.006 0.188 $\pm$ 0.002 0.215 $\pm$ 0.002 - 0.882 $\pm$ 0.006 -
BetaRCE(0.9) 0.455 $\pm$ 0.006 0.610 $\pm$ 0.006 0.224 $\pm$ 0.002 0.229 $\pm$ 0.002 - 0.926 $\pm$ 0.005 -
Seed BetaRCE(0.8) 0.308 $\pm$ 0.005 0.480 $\pm$ 0.006 0.177 $\pm$ 0.002 0.210 $\pm$ 0.002 - - 0.914 $\pm$ 0.005
BetaRCE(0.9) 0.399 $\pm$ 0.006 0.555 $\pm$ 0.006 0.205 $\pm$ 0.002 0.217 $\pm$ 0.002 - - 0.954 $\pm$ 0.004
Breast Cancer Standard CFEs Dice - 3.004 $\pm$ 0.003 1.157 $\pm$ 0.001 1.113 $\pm$ 0.001 0.894 $\pm$ 0.003 0.816 $\pm$ 0.003 0.838 $\pm$ 0.003
GrowingSpheres - 3.811 $\pm$ 0.001 0.864 $\pm$ 0.001 0.949 $\pm$ 0.001 0.898 $\pm$ 0.003 0.886 $\pm$ 0.003 0.886 $\pm$ 0.003
Robust post-hoc RobX (0.5,0.01) 1.251 $\pm$ 0.001 3.135 $\pm$ 0.001 0.706 $\pm$ 0.001 0.728 $\pm$ 0.001 0.848 $\pm$ 0.007 0.957 $\pm$ 0.004 0.846 $\pm$ 0.007
RobX (0.5,0.1) 3.163 $\pm$ 0.001 3.956 $\pm$ 0.001 0.876 $\pm$ 0.001 0.563 $\pm$ 0.001 0.997 $\pm$ 0.001 1.000 $\pm$ 0.001 0.996 $\pm$ 0.001
Arch BetaRCE(0.8) 1.684 $\pm$ 0.026 4.108 $\pm$ 0.04 0.916 $\pm$ 0.009 0.969 $\pm$ 0.007 0.925 $\pm$ 0.005 - -
BetaRCE(0.9) 2.058 $\pm$ 0.028 4.324 $\pm$ 0.042 0.964 $\pm$ 0.009 0.996 $\pm$ 0.007 0.961 $\pm$ 0.004 - -
Btsp BetaRCE(0.8) 1.528 $\pm$ 0.024 3.937 $\pm$ 0.041 0.901 $\pm$ 0.009 0.996 $\pm$ 0.008 - 0.926 $\pm$ 0.005 -
BetaRCE(0.9) 1.965 $\pm$ 0.026 4.176 $\pm$ 0.042 0.954 $\pm$ 0.01 1.024 $\pm$ 0.008 - 0.955 $\pm$ 0.004 -
Seed BetaRCE(0.8) 1.749 $\pm$ 0.028 3.788 $\pm$ 0.039 0.863 $\pm$ 0.009 0.940 $\pm$ 0.007 - - 0.920 $\pm$ 0.005
BetaRCE(0.9) 2.115 $\pm$ 0.032 4.000 $\pm$ 0.042 0.91 $\pm$ 0.009 0.967 $\pm$ 0.007 - - 0.945 $\pm$ 0.004
Car eval Standard CFEs Dice - 1.189 $\pm$ 0.001 $0.899\pm 0.001$ 0.469 $\pm$ 0.001 0.866 $\pm$ 0.003 0.800 $\pm$ 0.004 0.825 $\pm$ 0.004
GrowingSpheres - $0.965\pm 0.001$ $0.469\pm 0.001$ $0.487\pm 0.001$ $0.592\pm 0.006$ $0.566\pm 0.005$ $0.608\pm 0.005$
FACE - 0.977 $\pm$ 0.001 $0.633\pm 0.001$ 0.441 $\pm$ 0.001 0.761 $\pm$ 0.004 0.826 $\pm$ 0.004 0.766 $\pm$ 0.004
Robust end-to-end RBR - $0.822\pm 0.001$ $0.505\pm 0.001$ $0.460\pm 0.001$ $0.603\pm 0.001$ $0.649\pm 0.001$ $0.661\pm 0.001$
ROAR - $1.637\pm 0.001$ $0.739\pm 0.001$ $0.746\pm 0.001$ $0.277\pm 0.001$ $0.170\pm 0.001$ $0.329\pm 0.002$
Robust post-hoc RobX(0.5,0.01) 0.254 $\pm$ 0.002 1.110 $\pm$ 0.013 $0.573\pm 0.006$ 0.474 $\pm$ 0.001 0.844 $\pm$ 0.007 0.842 $\pm$ 0.007 0.903 $\pm$ 0.006
RobX(0.5,0.1) 0.895 $\pm$ 0.006 1.604 $\pm$ 0.015 $0.876\pm 0.007$ 0.456 $\pm$ 0.001 0.980 $\pm$ 0.003 0.995 $\pm$ 0.001 0.994 $\pm$ 0.001
BetaRCE Arch BetaRCE (0.8) 0.277 $\pm$ 0.014 0.948 $\pm$ 0.031 $0.456\pm 0.014$ 0.474 ±$\pm 0.002$ 0.817 $\pm$ 0.021 - -
BetaRCE (0.9) 0.353 $\pm$ 0.018 0.966 $\pm$ 0.027 $0.481\pm 0.013$ 0.474 ±$\pm 0.002$ 0.850 $\pm$ 0.020 - -
BetaRCE Btsr BetaRCE (0.8) 0.279 $\pm$ 0.005 1.148 $\pm$ 0.013 $0.561\pm 0.006$ 0.487 $\pm$ 0.001 - 0.925 $\pm$ 0.005 -
BetaRCE (0.9) 0.324 $\pm$ 0.006 1.172 $\pm$ 0.013 $0.573\pm 0.006$ 0.485 $\pm$ 0.001 - 0.948 $\pm$ 0.004 -
BetaRCE Seed BetaRCE (0.8) 0.27 $\pm$ 0.005 1.154 $\pm$ 0.013 $0.561\pm 0.006$ 0.487 $\pm$ 0.001 - - 0.939 $\pm$ 0.005
BetaRCE (0.9) 0.331 $\pm$ 0.005 1.193 $\pm$ 0.013 $0.581\pm 0.006$ 0.484 $\pm$ 0.001 - - 0.972 $\pm$ 0.003
Rice Standard CFEs Dice - 0.905 $\pm$ 0.001 $0.681\pm 0.001$ 0.493 $\pm$ 0.001 0.791 $\pm$ 0.004 0.804 $\pm$ 0.004 0.751 $\pm$ 0.004
GrowingSpheres - $0.863\pm 0.001$ $0.391\pm 0.001$ $0.250\pm 0.001$ $0.615\pm 0.005$ $0.669\pm 0.005$ $0.530\pm 0.005$
FACE - 0.805 $\pm$ 0.001 $0.341\pm 0.001$ 0.076 $\pm$ 0.001 0.619 $\pm$ 0.005 0.611 $\pm$ 0.005 0.566 $\pm$ 0.005
Robust end-to-end RBR - $0.890\pm 0.001$ $0.396\pm 0.001$ $0.213\pm 0.001$ $0.413\pm 0.001$ $0.421\pm 0.001$ $0.425\pm 0.001$
ROAR - $1.974\pm 0.001$ $0.813\pm 0.001$ $0.606\pm 0.001$ $0.231\pm 0.001$ $0.340\pm 0.002$ $0.284\pm 0.001$
Robust post-hoc RobX(0.5,0.01) 0.413 $\pm$ 0.004 1.036 $\pm$ 0.008 $0.445\pm 0.004$ 0.141 $\pm$ 0.002 0.940 $\pm$ 0.005 0.979 $\pm$ 0.003 0.953 $\pm$ 0.004
RobX(0.5,0.1) 1.021 $\pm$ 0.005 1.627 $\pm$ 0.009 $0.694\pm 0.004$ 0.101 $\pm$ 0.001 1.000 $\pm$ 0.001 1.000 $\pm$ 0.001 1.000 $\pm$ 0.001
BetaRCE Arch BetaRCE (0.8) 0.143 $\pm$ 0.003 1.063 $\pm$ 0.013 $0.487\pm 0.006$ 0.245 ±$\pm 0.001$ 0.821 $\pm$ 0.021 - -
BetaRCE (0.9) 0.205 $\pm$ 0.003 1.070 $\pm$ 0.013 $0.491\pm 0.006$ 0.250 ±$\pm 0.001$ 0.850 $\pm$ 0.020 - -
BetaRCE Btsr BetaRCE (0.8) 0.138 $\pm$ 0.003 1.167 $\pm$ 0.009 $0.591\pm 0.004$ 0.265 $\pm$ 0.001 - 0.895 $\pm$ 0.004 -
BetaRCE (0.9) 0.187 $\pm$ 0.004 1.172 $\pm$ 0.009 $0.593\pm 0.004$ 0.266 $\pm$ 0.001 - 0.923 $\pm$ 0.003 -
BetaRCE Seed BetaRCE (0.8) 0.286 $\pm$ 0.005 1.156 $\pm$ 0.009 $0.587\pm 0.004$ 0.261 $\pm$ 0.001 - - 0.919 $\pm$ 0.004
BetaRCE (0.9) 0.345 $\pm$ 0.005 1.162 $\pm$ 0.009 $0.589\pm 0.004$ 0.263 $\pm$ 0.001 - - 0.953 $\pm$ 0.003
Table 7: Comparative study results when the underlying model is a neural network and both RobX   and BetaRCE are using Dice as the base counterfactual explainer.